# AutoAgents: A Framework for Automatic Agent Generation

### v1

@@ 4.2 Trivia Creative Writing
Task Description. The Trivia Creative Writing task (Wang et al., 2023c) challenges the capabilities of large language models to retrieve and integrate diverse information from their internal self-compressed knowledge. This task requires a model to craft a coherent story around a given topic while incorporating the answers to $N$ trivia questions. We evaluate the models under two settings, $N=5$ and $N=10$, where a higher $N$ entails more trivia questions and thus demands the model to exhibit more extensive domain knowledge. We constructed a benchmark consisting of 100 instances for each $N$, encompassing a total of 1000 trivia questions.
@@
Methods N (# trivia questions) = 5 N (# trivia questions ) = 10 Score (%) $\Delta$ (v.s Standard %) Score (%) $\Delta$ (v.s Standard %) Standard 74.6 0.0% 77.0 0.0% CoT (Yao et al., 2023) 67.1 -10.0% 68.5 -11.1% SPP-Profile (Wang et al., 2023c) 79.1 +5.9% 83.0 +7.8% SPP (Wang et al., 2023c) 79.9 +7.1% 84.7 +10.0% AutoAgents 82.0 +9.9% 85.3 +10.8%
@@
Results. Table 3 demonstrates the superior performance of AutoAgents in knowledge acquisition over the existing methods. Compared to the Standard method, which does not employ Agent Generation, AutoAgents achieves a remarkable 10% improvement across all experiments. Moreover, AutoAgents also surpasses SSP (Wang et al., 2023c), which utilizes agent generation but with a different approach. The enhanced performance of AutoAgents can be attributed to its elaborate methods of agent generation discussions and task execution including collaborative refinement and self-refinement. More examples are given in the Appendix 4.4.
@@ 4.3 Case Study
4.4 Further Analysis
Collaborative discussion is crucial for rational agent generation and plan allocation. During the Drafting Stage, the Planner in AutoAgents engages in collaborative discussions with two Observers to determine the optimal list of agents and the execution plan. Figure 5 illustrates the contrast between agent generation with and without collaborative discussion. In the absence of Observer feedback, the Planner tends to generate programmers exclusively to accomplish game development, neglecting the holistic process of game creation. With the input and coordination of the Observers, the Planner incorporates game design experts, UI design experts, and testing experts into the agent list. It is evident that the agent generation under collaborative discussions is more comprehensive and more aligned with the realistic scenarios of game development. This also corroborates the significance of collaborative discussions for agent generation and plan allocation, which will subsequently influence the execution outcomes.
Enhancing single-agent and multi-agent collaboration through refinement. Self-Refinement (Madaan et al., 2023; Shinn et al., 2023; Gou et al., 2023; Chen et al., 2023b; Huang et al., 2022; Yao et al., 2022a) is a technique that enables LLMs to “converse” with themselves, evaluate their own generation, and iteratively improve their answers. Self-refinement has been shown to enhance the accuracy of LLMs’ outputs in various domains (Madaan et al., 2023; Shinn et al., 2023; Gou et al., 2023; Chen et al., 2023b; Huang et al., 2022; Yao et al., 2022a). Although AutoAgents is a framework for multi-agent collaboration, it also requires self-refinement agents to perform specialized roles for individual tasks. Figure 1 depicts the self-refinement process of programmers’ coding. They first write a pseudo code file and then generate the corresponding program files based on it. This refinement process significantly ensures the validity of the output file. For collaborative refinement, the process resembles the collaborative dialogue mentioned above, which involves integrating knowledge from different domains to accomplish tasks that demand cross-domain knowledge fusion. In summary, refinement is vital for both self-correction and collaborative communication.
@@ 4.4 Further Analysis
Figure 7: An example of the Open-ended Question Answer.
Dynamic agents enhance the adaptability of complex tasks. The ability to generate dynamic agents for various tasks is crucial for enhancing their adaptability to diverse scenarios. Figure 7 illustrates the contrast between GPT4 and AutoAgents’ responses to open-ended questions. Unlike GPT-4, AutoAgents can produce agents from three distinct domains, which can provide more elaborate answers. For the trivia creative writing task in Figure 8 and 9, AutoAgents decomposed the task into four steps: first, it retrieved the answer to the question through a domain-specific agent and then constructed the story. Meanwhile, it is evident that the Language Expert Agent performed multiple checks on the coherence between the story content and the question, ensuring the accuracy of the story content. These examples all demonstrate the versatility of building a dynamic agent generation framework for complex tasks.
@@
5 Discussion
@@
Future Work. Cooperation among multiple agents requires dynamic adaptation and communication. The initial plan generated by LLMs may not suffice to achieve the desired outcomes, resulting in erroneous final output results. Hence, future multi-agent systems need to swiftly detect and rectify errors and adjust their plans dynamically to align with the desired outcomes.

### v3

@@ 4.2 Trivia Creative Writing
Task Description. The Trivia Creative Writing task [34] challenges the capabilities of large language models to retrieve and integrate diverse information from their internal self-compressed knowledge. This task requires a model to craft a coherent story around a given topic while incorporating the answers to $N$ trivia questions. We evaluate the models under two settings, $N=5$ and $N=10$, where a higher $N$ entails more trivia questions and thus demands the model to exhibit more extensive domain knowledge. We constructed a benchmark consisting of 100 instances for each $N$, encompassing a total of 1000 trivia questions.
@@
Methods N (# trivia questions) = 5 N (# trivia questions ) = 10 Score (%) $\Delta$ (v.s Standard %) Score (%) $\Delta$ (v.s Standard %) Standard 74.6 0.0% 77.0 0.0% CoT [39] 67.1 -10.0% 68.5 -11.1% SPP-Profile [34] 79.1 +5.9% 83.0 +7.8% SPP [34] 79.9 +7.1% 84.7 +10.0% AutoAgents 82.0 +9.9% 85.3 +10.8%
@@
Results. Table 3 demonstrates the superior performance of AutoAgents in knowledge acquisition over the existing methods. Compared to the Standard method, which does not employ Agent Generation, AutoAgents achieves a remarkable 10% improvement across all experiments. Moreover, AutoAgents also surpasses SSP [34], which utilizes agent generation but with a different approach. The enhanced performance of AutoAgents can be attributed to its elaborate methods of agent generation discussions and task execution including collaborative refinement and self-refinement. More examples are given in the Appendix A.
4.3 Further Analysis
This section delves into the significance of key components within AutoAgents by separately analyzing the self-refinement action, collaborative refinement action, dynamic memory, and observers in the draft stage across 20 instances  [^3] The last 20 samples from a dataset of 100 samples are used as test instances. of the Trivia Creative Writing task and additional case studies.
Table 4: The ablation studies of AutoAgents on 20 instances of Trivia Creative Writing task. $\Delta$ indicates the differences compared with Standard Prompting (first row).
Methods N (# trivia questions) = 5 Score (%) $\Delta$ (v.s Standard %) Standard 74.6 0.0% CoT [39] 66.0 -11.5% SPP-Profile [34] 74.0 -0.01% SPP [34] 84.4 +13.1% AutoAgents w/o observers 87.0 +16.6% AutoAgents w/o self-refinement 87.0 +16.6% AutoAgents w/o collaborative refinement 88.0 +18.0% AutoAgents w/o dynamic memory 89.0 +19.3% AutoAgents 90.0 +20.6%
Collaborative discussion is crucial for rational agent generation and plan allocation. During the Drafting Stage, the Planner in AutoAgents engages in collaborative discussions with two Observers to determine the optimal list of agents and the execution plan. Figure 5 illustrates the contrast between agent generation with and without collaborative discussion. In the absence of Observer feedback, the Planner tends to generate programmers exclusively to accomplish game development, neglecting the holistic process of game creation. With the input and coordination of the Observers, the Planner incorporates game design experts, UI design experts, and testing experts into the agent list. It is evident that the agent generation under collaborative discussions is more comprehensive and more aligned with the realistic scenarios of game development. This also corroborates the significance of collaborative discussions for agent generation and plan allocation, which will subsequently influence the execution outcomes. Concurrently, Table 4 elucidates that in the absence of observers, there is a marked 3% reduction in the overall performance of AutoAgents. This substantiates the imperative role of collaborative discussions in agent generation. AutoAgent markedly enhances the caliber of agent generation via collaborative discussions, a facet notably overlooked by other generative frameworks in their consideration of agent generation quality. The empirical data presented in Table 2 and  3 further accentuate the superiority of AutoAgents when juxtaposed against counterparts like AgentVerse and SPP.
Enhancing single-agent through self-refinement. Self-Refinement [18, 28, 9, 6, 13, 40] is a technique that enables LLMs to “converse” with themselves, evaluate their own generation, and iteratively improve their answers. Self-refinement has been shown to enhance the accuracy of LLMs’ outputs in various domains [18, 28, 9, 6, 13, 40]. Although AutoAgents is a framework for multi-agent collaboration, it also requires self-refinement agents to perform specialized roles for individual tasks. As shown in the results in Table 4, the performance of AutoAgents decreases by 3% in the absence of the self-refinement action. This observation corroborates the assertion that self-refinement is instrumental in augmenting proficiency in trivia creative writing tasks. Furthermore, the enhancement of single agents via self-refinement plays a pivotal role in fortifying the integrity of the overarching multi-agent framework.
Enhancing multi-agent collaboration through collaborative refinement action. For collaborative refinement, the process resembles the collaborative dialogue mentioned above, which involves integrating knowledge from different domains to accomplish tasks that demand cross-domain knowledge fusion. The results in Table 4 demonstrate the performance when collaborative refinement is absent. It’s observable that compared to the scenario with AutoAgents, there is a decline of 2%. Since the necessity for multiple agents to collaborate on a single task is entirely dependent on the decision made by the agent in the drafting phase, not all problems necessarily involve tasks that require collaborative refinement. However, it’s evident that when this principle is omitted from the prompt’s design, there’s a noticeable performance decrease.
Improve the effectiveness of actions by dynamic memory. Dynamic memory predominantly addresses the requisites of specialized agents. As shown in Figure 4, the Action Observer amalgamates pivotal data for forthcoming tasks, utilizing the historical action records archived in long-term memory. Table 4 elucidates a 1% diminution in the efficacy of AutoAgents bereft of dynamic memory. Quintessential insights derived from dynamic memory are assimilated into the prompt, thereby augmenting the comprehension of critical information and bolstering the operational proficiency of actions.
@@ 4.4 Case Study
5 Conclusion
This paper introduces AutoAgents, an innovative framework for automatically synthesizing collaborative specialized agents. AutoAgents mimics the collaborative process of human teams by decomposing tasks into drafting and execution phases and delegating different subtasks to different agents. Our experimental and empirical evaluation validates the advantages of AutoAgents, as it surpasses single agents and other groupings in various tasks that demand diverse skills. Furthermore, our case study in software development illustrates the versatility and potential benefits of our proposed framework. AutoAgents opens up new possibilities for enhancing the interaction and cooperation among agents and transforms the landscape of complex problem-solving. We envisage that its principles can be further generalized and refined to deal with a broader range of tasks, paving the way towards more useful assistive AI.
6 Acknowledgements
This work was partially supported by the National Key R&D Program of China (2022YFC2009600 and 2022YFC2009606) and the Postdoctoral Fellowship Program of CPSF under Grant Number GZB20230024.
Appendix A More Examples
Enhancing single-agent through self-refinement. Figure 7 depicts the self-refinement process of programmers’ coding. They first write a pseudo code file and then generate the corresponding program files based on it. This refinement process significantly ensures the validity of the output file. Although AutoAgents is a framework for multi-agent collaboration, it also requires self-refinement agents to perform specialized roles for individual tasks.
Figure 7: An example of the self-refinement process of programmers’ coding
Improving Teamwork in Multi-Agent Systems with Collaborative Refinement. In collaborative refinement, the method is similar to the team discussions we talked about earlier. This involves bringing together information from various fields to complete tasks that require blending knowledge from different areas. As illustrated in Figure 8, during this process, the two agents work together to make sure the story they produce makes sense and is consistently improved upon. This also supports the idea that working together in this way is helpful when dealing with complicated tasks.
Figure 8: An example of the collaborative refinement process of trivia creative writing.
@@ Appendix A More Examples
Dynamic agents enhance the adaptability of complex tasks. The ability to generate dynamic agents for various tasks is crucial for enhancing their adaptability to diverse scenarios. Figure 9 illustrates the contrast between GPT4 and AutoAgents’ responses to open-ended questions. Unlike GPT-4, AutoAgents can produce agents from three distinct domains, which can provide more elaborate answers. For the trivia creative writing task in Figure 10 and 11, AutoAgents employs a four-step approach for task decomposition. Initially, it sources the answer to the given question using a domain-specific agent, followed by the construction of a narrative. Concurrently, the Language Expert Agent plays a pivotal role, conducting multiple checks to verify the coherence between the narrative and the question, thus guaranteeing the narrative’s accuracy.
@@
Furthermore, as illustrated in Figure 12, the Action Observer orchestrates the interaction among multiple generative agents. It provides a concise summary of essential information, proving instrumental in fostering collaboration between various intelligent agents. This coordination is key to ensuring the seamless flow of the task execution process. Collectively, these instances vividly showcase the adaptability and efficiency of our dynamic agent generation framework in handling complex tasks.
Figure 12: An example of the coordination process for the Acton Observer agent.
Table 5: Comparison of existing and proposed frameworks for multi-agent generation methods.
Framework Application Agent Generalization by Multi-Agent Discussion Prompt Generalization Social Simulacra [24] Social Simulation Epidemic Modeling [35] Social Simulation SSP [34] General Autonomous Agents AgentVerse [5] General Autonomous Agents AutoAgents General Autonomous Agents
Conversely, the prompt employed by AutoAgents exhibits a more universal nature, signifying its capacity to acclimate to diverse tasks without necessitating bespoke customization. As Table 5 delineates, both AgentVerse  [^4] https://github.com/OpenBMB/AgentVerse/tree/minecraft/agentverse/tasks and SSP  [^5] https://github.com/MikeWangWZHL/Solo-Performance-Prompting/tree/main/prompts have implemented task-specific enhancements for varied task evaluations. In contrast, our methodology leverages a singular, unified prompt format to accommodate an array of tasks. The commendable efficacy in open-ended question-answer and trivia creative writing tasks further corroborates the wide-ranging applicability and versatility of prompt design within the AutoAgents framework.
Appendix B Human Evaluation
In this section, we present the criteria for human evaluation. We instructed the volunteers, who are responsible for assessing the quality of different feedback, to adhere to these standards.
[Text]Human Evaluation
We would like to request your feedback on the response to the user questiondisplayed above. Please rate the helpfulness, relevance, accuracy, level ofdetails of their responses.Each response receives an overall score on a scale of 1 to 10, where a higherscore indicates better overall performance. Please first provide acomprehensive explanation of your evaluation, avoiding any potential bias andensuring that the order in which the responses were presented does not affectyour judgment.Output with the following format:Evaluation evidence: <your evluation explanation here>Score: <score>
Appendix C Discussion
@@
Future Work. Cooperation among multiple agents requires dynamic adaptation and communication. The initial plan generated by LLMs may not suffice to achieve the desired outcomes [4], resulting in erroneous final output results. Hence, future multi-agent systems need to swiftly detect and rectify errors and adjust their plans dynamically to align with the desired outcomes.